OpenStudy (anonymous):

lim_{x rightarrow 1}sqrt{x+3}-2over lim_{x rightarrow 1}sqrt{10-x}-3

OpenStudy (anonymous):

\[\lim_{x \rightarrow 1}\sqrt{x+3}-2\over \lim_{x \rightarrow 1}\sqrt{10-x}-3\]

OpenStudy (anonymous):

0 \or 0/0

OpenStudy (anonymous):

no thats not the answer

OpenStudy (anonymous):

-1 final answer

OpenStudy (anonymous):

how do you get that?

OpenStudy (dumbcow):

when you initially evaluate a lim and get an indeterminate answer such as 0/0 use L'hopitals rule which says an equivalent limit can be obtained by differentiating the top and bottom then re-evaluating After differentiating: \[\lim_{x \rightarrow 1} \frac{\frac{1}{2\sqrt{x+3}}}{-\frac{1}{2\sqrt{10-x}}} = \lim_{x \rightarrow 1} -\frac{\sqrt{10-x}}{\sqrt{x+3}} = -\frac{3}{2}\]